OpenStudy (anonymous):

13/11 ~greater than, less than or equal to~ 8/7??

OpenStudy (anonymous):

same principle as the last problem just multiply by the denominator of the other one so 13/11 * 7/7 = 91/77 8/7 * 11/11 = 88/77